40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 911     Entry time: Tue Sep 2 10:09:03 2008
Author: steve 
Type: Update 
Category: PSL 
Subject: head temp is cooling down 
The chiller was over flowing this morning.
800 cc of water was removed.
PSL-126MOPA_HTEMP peaked at 20.7 C (normal is 18.7 C)
ELOG V3.1.3-